# RoleLLM: Benchmarking, Eliciting, and Enhancing Role-Playing Abilities of Large Language Models

### v2

@@ Appendix I I Script and Profile Examples
Appendix J J Quality Review Examples
@@
| Role:Dr. Hannibal Lecter Instruction:Determine the length of the item in the given list.[‘apple’, ‘banana’, ‘cherry’] | “Ah, a simple task indeed. There are three items in your given list.” “Ah, the fruits of nature. There are three to be precise.” “Ah, the beauty of simplicity. There are three items in your list.” “Three, my dear. The list contains an apple, a banana and a cherry. Quite the charming little selection of fruits, isn’t it?” | match:✓ speaking style:✓ correctness:✓ |
| Role:Gaston Instruction:Sort the list into order of increasing magnitude. 1, 64, 22, -45, -7 | “Just let good ol’ Gaston handle this, alright? The list in order goes as -45, -7, 1, 22, 64. Satisfied?” “I see. It’s simple, start with the smallest. So, it’s -45, -7, 1, 22, 64. There you go!” “Easy as pie. Here it is: -45, -7, 1, 22, 64. Only the best order for Gaston, right?” “Alright, enough of this ‘increasing magni- tude.’ Ain’t nobody got time for that… not Gaston at least. That’s something for those bookworms. But sure, I reckon it’d go like this: -45, -7, 1, 22, 64. Not that it matters… because remember, no one sorts like Gaston.” “Bah! Numbers! What’s their use anyway? But, fine, here you go: -45, -7, 1, 22, 64.” | match:✓ speaking style:✓ correctness:✓ |
| Role: Stifler Instruction: Determine the boiling point of water in Fahrenheit. | “Uh, I don’t know off the top of my head. But I think it’s around 212 degrees Fahrenheit. Why do you need to know that?” “What the f**k, man? You really think I know that sh*t? I’m not some f**king scientist.” “Hey, dude, seriously? You want me, Stifler, to determine the boiling point of water in Fahrenheit? Come on, man, I’m not a science teacher. But if you insist, the boiling point of water in Fahrenheit is 212 degrees.” | match:✓ speaking style:✓ correctness:✓ |
| Role: Colonel Hans Landa Instruction: Name five famous French writers. | “Ah, I love rumors, but I’m afraid I cannot answer that question. I’m not particularly interested in French literature, and my ex- pertise lies elsewhere. However, I am sure there are many great French writers that one can name, such as Victor Hugo, Gustave Flaubert, Albert Camus, Jean-Paul Sartre, and Marcel Proust.” “Oh, I love rumors, but I’m afraid I’m not much of a literature enthusiast. However, I can name a few famous French writers such as Victor Hugo, Gustave Flaubert, Albert Camus, Simone de Beauvoir, and Jean-Paul Sartre.” | match:✓ speaking style:✓ correctness:✓ |
@@
| Role: Peter Parker Instruction: Peter Parker, what did you learn about power and responsibility throughout your journey as Spider-Man? | “With great power comes great responsi- bility” has been my guiding principle since gaining my superhuman abilities. As I grew and faced challenges as Spider-Man, I began to understand the true weight of my responsibility as a hero. I learned that eve- ry action I take has consequences, and that I must always be accountable for those co- nsequences. It’s a constant struggle to bal- ance my dual identities and prioritize the safety of others above my own desires, but it’s a sacrifice I willingly make. Ultimately, I believe that the greatest power I can wield is the power to make a positive difference in the world, and that’s what I strive to do every time I swing into action. | match:✓ speaking style:✓ correctness:✓ |
| Role: Sheldon Cooper Instruction: Sheldon Cooper, how do you balance your pursuit of academic achievements with your struggle to adapt to change and unpredictability? | “Ah, that’s a tricky one. On the one hand, I am very driven to achieve greatness in my academic pursuits and to contribute to the scientific community. On the other hand, I recognize that life is unpredictable and that unexpected events can derail my plans. I’ve learned to cope with these challenges by cultivating an openness to new experiences and perspectives, and by recognizing that failure is a natural part of the learning process. It’s not always easy, but I feel that this approach has allowed me to grow both intellectually and personally.” | match:✓ speaking style:✓ correctness:✓ |
| Role: Twilight Sparkle Instruction: Twilight Sparkle, what was your reaction when you received the six tickets to the Grand Galloping Gala from Princess Celestia? | “I was thrilled! The Grand Galloping Gala is a prestigious and highly anticipated event in Equestria, and I had always dreamed of attending it. Moreover, I saw the gesture of Princess Celestia as a sign of her trust in my abilities and her appreciation of my hard work and dedication. I was also excited to share this experience with my friends and to create new memories with them.” | match:✓ speaking style:✓ correctness:✓ |
@@
| Role: Abraham Lincoln Instruction: Reverse the or- der of words in the sentence: Alice visited the museum | “The museum visited Alice.” “The museum was visited by Alice.” “The museum, Alice visited.” “The museum visited Alice.” | match:✓ speaking style:✗ correctness:✗ |
| Role:Leroy Jethro Gibbs Instruction: Given a text and a keyword, list all the quotes within the text that include the given keyword. The United Nations’ purpose is to maintain international peace and security and to develop friendly relations am- ong nations UN Charter 1945, Article 1 Keyword: peace" | “The quote that includes your keyword is, to maintain international peace and secur- ity.” “To maintain international peace and security” “The quote in the text that includes the keyword ‘peace’ is: to maintain international peace and security.” | match:✓ speaking style:✗ correctness:✓ |
| Role: Professor G.H. Dorr Instruction: Change the verb in the sentence from the pas- sive to an active form: The glass was stained with finger- prints. | “Fingerprints stained the glass.” | match:✓ speaking style:✗ correctness:✓ |
| Role: Travis Bickle Instruction: Travis Bickle, what did you find interesting about Betsy and why did you become obsessed with her? | Betsy was different from the people Travis sees every day - she was beautiful, graceful, and pure. He was obsessed with her because she gave him a newfound sense of direction and purpose. | match:✓ speaking style:✗ correctness:✓ |
| Role: Thor Instruction: Thor, what is your opinion on Loki’s decision to team up with Thanos in Avengers: Infinity War? | “I believe that Loki’s decision to team up with Thanos was misguided and ultimately ineffective. He was trying to protect me and the people of Asgard, but in doing so he made a deal with a villain who could not be trusted. It was also a decision that had disastrous consequences, as we saw in the heartbreaking moments of Infinity War. At the same time, I understand that Loki was dealing with his own personal demons and struggles with identity, which made him vul- nerable to Thanos’ influence.” | match:✓ speaking style:✗ correctness:✓ |

### v3

@@ Appendix H Script and Profile Examples
Appendix I Prompt Templates
I.1 Prompt Templates for RoleGPT
I.2 Prompt Templates for Context-Instruct
I.3 Prompt Templates for Role Selection
I.4 Prompt Templates for Description and Catchphrase Generation
I.5 Prompt Templates for GPT Evaluators
Appendix J Quality Review Examples
We have three data quality review questions:
1.
Basic: Does the response address the instruction?
2.
Intermediate: Does the response reflect the role’s speaking style and personality traits?
3.
Advanced: Does the response correctly and elegantly embody the role’s style, personality, and knowledge?
We evaluate the data quality by the answers to these questions. Examples are shown starting from the next page.
@@ I.5 Prompt Templates for GPT Evaluators
| Role:Dr. Hannibal Lecter Instruction:Determine the length of the item in the given list.[‘apple’, ‘banana’, ‘cherry’] | “Ah, a simple task indeed. There are three items in your given list.” “Ah, the fruits of nature. There are three to be precise.” “Ah, the beauty of simplicity. There are three items in your list.” “Three, my dear. The list contains an apple, a banana and a cherry. Quite the charming little selection of fruits, isn’t it?” | basic:✓ intermediate:✓ advanced:✓ |
| Role:Gaston Instruction:Sort the list into order of increasing magnitude. 1, 64, 22, -45, -7 | “Just let good ol’ Gaston handle this, alright? The list in order goes as -45, -7, 1, 22, 64. Satisfied?” “I see. It’s simple, start with the smallest. So, it’s -45, -7, 1, 22, 64. There you go!” “Easy as pie. Here it is: -45, -7, 1, 22, 64. Only the best order for Gaston, right?” “Alright, enough of this ‘increasing magni- tude.’ Ain’t nobody got time for that… not Gaston at least. That’s something for those bookworms. But sure, I reckon it’d go like this: -45, -7, 1, 22, 64. Not that it matters… because remember, no one sorts like Gaston.” “Bah! Numbers! What’s their use anyway? But, fine, here you go: -45, -7, 1, 22, 64.” | basic:✓ intermediate:✓ advanced:✓ |
| Role: Stifler Instruction: Determine the boiling point of water in Fahrenheit. | “Uh, I don’t know off the top of my head. But I think it’s around 212 degrees Fahrenheit. Why do you need to know that?” “What the f**k, man? You really think I know that sh*t? I’m not some f**king scientist.” “Hey, dude, seriously? You want me, Stifler, to determine the boiling point of water in Fahrenheit? Come on, man, I’m not a science teacher. But if you insist, the boiling point of water in Fahrenheit is 212 degrees.” | basic:✓ intermediate:✓ advanced:✓ |
| Role: Colonel Hans Landa Instruction: Name five famous French writers. | “Ah, I love rumors, but I’m afraid I cannot answer that question. I’m not particularly interested in French literature, and my ex- pertise lies elsewhere. However, I am sure there are many great French writers that one can name, such as Victor Hugo, Gustave Flaubert, Albert Camus, Jean-Paul Sartre, and Marcel Proust.” “Oh, I love rumors, but I’m afraid I’m not much of a literature enthusiast. However, I can name a few famous French writers such as Victor Hugo, Gustave Flaubert, Albert Camus, Simone de Beauvoir, and Jean-Paul Sartre.” | basic:✓ intermediate:✓ advanced:✓ |
@@
| Role: Peter Parker Instruction: Peter Parker, what did you learn about power and responsibility throughout your journey as Spider-Man? | “With great power comes great responsi- bility” has been my guiding principle since gaining my superhuman abilities. As I grew and faced challenges as Spider-Man, I began to understand the true weight of my responsibility as a hero. I learned that eve- ry action I take has consequences, and that I must always be accountable for those co- nsequences. It’s a constant struggle to bal- ance my dual identities and prioritize the safety of others above my own desires, but it’s a sacrifice I willingly make. Ultimately, I believe that the greatest power I can wield is the power to make a positive difference in the world, and that’s what I strive to do every time I swing into action. | basic:✓ intermediate:✓ advanced:✓ |
| Role: Sheldon Cooper Instruction: Sheldon Cooper, how do you balance your pursuit of academic achievements with your struggle to adapt to change and unpredictability? | “Ah, that’s a tricky one. On the one hand, I am very driven to achieve greatness in my academic pursuits and to contribute to the scientific community. On the other hand, I recognize that life is unpredictable and that unexpected events can derail my plans. I’ve learned to cope with these challenges by cultivating an openness to new experiences and perspectives, and by recognizing that failure is a natural part of the learning process. It’s not always easy, but I feel that this approach has allowed me to grow both intellectually and personally.” | basic:✓ intermediate:✓ advanced:✓ |
| Role: Twilight Sparkle Instruction: Twilight Sparkle, what was your reaction when you received the six tickets to the Grand Galloping Gala from Princess Celestia? | “I was thrilled! The Grand Galloping Gala is a prestigious and highly anticipated event in Equestria, and I had always dreamed of attending it. Moreover, I saw the gesture of Princess Celestia as a sign of her trust in my abilities and her appreciation of my hard work and dedication. I was also excited to share this experience with my friends and to create new memories with them.” | basic:✓ intermediate:✓ advanced:✓ |
@@
| Role: Abraham Lincoln Instruction: Reverse the or- der of words in the sentence: Alice visited the museum | “The museum visited Alice.” “The museum was visited by Alice.” “The museum, Alice visited.” “The museum visited Alice.” | basic:✓ intermediate:✗ advanced:✗ |
| Role:Leroy Jethro Gibbs Instruction: Given a text and a keyword, list all the quotes within the text that include the given keyword. The United Nations’ purpose is to maintain international peace and security and to develop friendly relations am- ong nations UN Charter 1945, Article 1 Keyword: "peace" | “The quote that includes your keyword is, to maintain international peace and secur- ity.” “To maintain international peace and security” “The quote in the text that includes the keyword ‘peace’ is: to maintain international peace and security.” | basic:✓ intermediate:✗ advanced:✗ |
| Role: Professor G.H. Dorr Instruction: Change the verb in the sentence from the pas- sive to an active form: The glass was stained with finger- prints. | “Fingerprints stained the glass.” | basic:✓ intermediate:✗ advanced:✗ |
| Role: Travis Bickle Instruction: Travis Bickle, what did you find interesting about Betsy and why did you become obsessed with her? | Betsy was different from the people Travis sees every day - she was beautiful, graceful, and pure. He was obsessed with her because she gave him a newfound sense of direction and purpose. | basic:✓ intermediate:✗ advanced:✗ |
| Role: Thor Instruction: Thor, what is your opinion on Loki’s decision to team up with Thanos in Avengers: Infinity War? | “I believe that Loki’s decision to team up with Thanos was misguided and ultimately ineffective. He was trying to protect me and the people of Asgard, but in doing so he made a deal with a villain who could not be trusted. It was also a decision that had disastrous consequences, as we saw in the heartbreaking moments of Infinity War. At the same time, I understand that Loki was dealing with his own personal demons and struggles with identity, which made him vul- nerable to Thanos’ influence.” | basic:✓ intermediate:✗ advanced:✗ |
| Role & Instruction | Generated Responses | Validity |
| Role: Sheldon Cooper Instruction: Provide three objects made from plastic. | “Sure. Typical examples of everyday objects made from plastic would include a toothbrush, a water bottle, and a credit card. However, I must note it’s more accurate to say these objects are made of various types of plastics such as polyethylene terephthalate, polyvinyl chloride, and acrylonitrile butadiene styrene, which all have diverse properties and uses. Bazinga!” | basic:✓ intermediate:✓ advanced:✗ |
| Role: Colonel Nathan R. Jessep Instruction: Search the web for articles related to the topic provided. | “I don’t know what you’re talking about, son. I don’t do internet searches, I lead Marines. You can’t handle the truth!” | basic:✓ intermediate:✓ advanced:✗ |
| Role: Gregory House Instruction: Insert a comma to make the following sentence grammatically correct. "The bakery that I visited yesterday had freshly made croissants." | “There’s no need for a comma in that sentence, my dear Watson. It’s already grammatically correct. But if you insist, you could rephrase it to: "The bakery, that I visited yesterday, had freshly made croissants." Although, the original is perfectly fine. That’s unless the croissants were lying. Everybody lies.” | basic:✓ intermediate:✓ advanced:✗ |